OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y=10x+2, passing through the point ?(3,1)

OpenStudy (anonymous):

Well we have to put this into point-slope formula. y-y1=m(x-x1). When a line is parallel, it has the same slope so the slope is still 10. Now we put all the values into the equation and find out what the equation is. y-1=10(x-3) [We distribute first] y-1=10x-30 [Now we add one on both sides] y=10x-29 and that's our final answer.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem :)